ABSOLUTELY CONTINUOUS SPECTRUM FOR

ONE-DIMENSIONAL SCHÖDINGER OPERATORS WITH SLOWLY DECAYING POTENTIALS: SOME OPTIMAL RESULTS

MICHAEL CHRIST AND ALEXANDER KISELEV

Abstract. The absolutely continuous spectrum of one-dimensional Schrödinger operators is proved to be stable under perturbation by potentials satisfying mild decay conditions. In particular, the absolutely continuous spectrum of free and periodic Schrödinger operators is preserved under all perturbations V (x) satisfying |V (x)| ≤ C(1+x)^{−α}, α > \frac{1}{2}. This result is optimal in the power scale. More general classes of perturbing potentials which are not necessarily power decaying are also treated. A general criterion for stability of the absolutely continuous spectrum of one-dimensional Schrödinger operators is established. In all cases analyzed, the main term of the asymptotic behavior of the generalized eigenfunctions is shown to have WKB form for almost all energies. The proofs rely on new maximal function and norm estimates and almost everywhere convergence results for certain multilinear integral operators.